OpenStudy (anonymous):

help with trig integration ? Stuck on 1 step lol

OpenStudy (anonymous):

\[\int\limits \sqrt{x^2-9}/(x)\]

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

i got stuck on integral 3tantheta *secthetatantheta/3sectheta

OpenStudy (amistre64):

if u = sqrt(x^2-9) du = 2x/2sqrt(x^2-9) dx seems like things simplify to me

OpenStudy (anonymous):

this is using trig sub

OpenStudy (amistre64):

\[\int\limits \sqrt{x^2-9}/(x)~dx\] \[\int\limits u~2\sqrt{x^2-9}/(x^2)~dx\] \[\int\limits 2u^2/(\sqrt u-9)~dx\] what trig sub did you try?

OpenStudy (anonymous):

using the chain rule d/du=sin((e^u)(u^2))=dsinu/du where u=e^u(u^2)

OpenStudy (anonymous):

x=3secTHETA dx = 3secTHETAtanTHETA

OpenStudy (amistre64):

sin^2 - 1 = -cos^2 cos^2 - 1 = -sin^2 tan^2 = sec^2 - 1 all come to mind for the -1 part

OpenStudy (anonymous):

Omg I just realized i made a dumb retricemistake thanks anyways!!

OpenStudy (amistre64):

yeah, it does tend to be one of those mistakes :) i tend to make more of them on a test lol

OpenStudy (anonymous):

Lmao same and then i just blank out after -_-

TheSmartOne (thesmartone):

Welcome to Open Study!!

OpenStudy (anonymous):

\[\sqrt{16-4\sin^2x}\]